OpenStudy (anonymous):

Point A is located at (-2, -6), and D is located at (-6, 8). Find the coordinates of the point that lies halfway between A and D. Select one: a. (0, 2) b. (-2, 3) c. (-3, 0) d. (-4, 1)

OpenStudy (anonymous):

@Abmon98 :)))

OpenStudy (nirmalnema):

use mid point formula.. (x1+x2) /2 (y1+y2)/2

OpenStudy (nirmalnema):

-2-6/ 2=-4=x -6+8/ 2 = 1=y

OpenStudy (abmon98):

(-2+(-6)/2,(-6+8/2)=(-4,1)

OpenStudy (nirmalnema):

option D (-4,1)